Please help me solve the equation and check the solution. Thanks!
17-7x+3 = -3x+21-3x

Answer by 10301650(32)

17-7x+3=-321-3x
-7x+3x+3x=21-3-17
-x=21-20
-x=1
-1(-x=1)-1
x=-1 (ANSWER)